A cake retailer increases the price of its cakes from £2.00 to £2.60 per cake and the quantity demanded decreases from 60 cakes per day to 45 cakes per day.

The price elasticity of demand for the retailer’s cake is


1
Expert's answer
2022-03-08T09:49:17-0500

"PED=\frac{\Delta Q}{\Delta P}×\frac{P}{Q}"

"=\frac{45-60}{2.6-2}×\frac{2}{60}"

"=\frac{-15}{0.6}×\frac{2}{60}= \frac{-30}{36}" = 0.833